OpenStudy (anonymous):

What is the rule for the sequence 1, 3, 9, 27, …? Answer Add 1 to the previous term Multiply the previous term by 2 Add 2 to the previous term Multiply the previous term by 3

OpenStudy (anonymous):

multiply previous term by 3

OpenStudy (unklerhaukus):

powers of three

Parth (parthkohli):

Powers of three starting from 0 ^

Parth (parthkohli):

\[\{3^n \}_{0\le n\le \infty} \]Fancy, fancy stuff.